OpenStudy (anonymous):

How do you find the domain and range for f(x)= x^2+x-2/x^2-3x-4 when factored it looks like this (x+2)(x-1)/(x+1)(x-4)

hero (hero):

I responded to your other question bro

OpenStudy (anonymous):

i know this is my last one i have a problem with.. im srry im just confused

hero (hero):

There are a couple of things to know which is 1. For the domain: The denominator cannot equal zero, therefore (x+1)(x-4) = 0 is invalid. In other words, the x values that make that expression true are invalid. 2. For the range: You need to find a value (if any) that would cause you not to be able to find an x value. If you try f(x) = 0, you'll just end up with 0 = (x+2)(x - 1) but that expression is in the numerator so those x values are valid. Now you'll have to think of another value for f(x) that would cause your function to be invalid. Can you think of anything? I can't.

OpenStudy (anonymous):

No i can't.. did i write it wrong? or is there no domain or range

hero (hero):

There is a domain and range. What I've done is help you figure out what values we can possibly eliminate from the domain or range. Basically, what you do with domain and range is assume that the domain is all real numbers and assume that the range is all real numbers unless you can find values that can be eliminated from the domain or range. Then the domain or range will be all real numbers EXCLUDING the values that are invalid.

OpenStudy (anonymous):

So the domain is -2 and 1?

hero (hero):

If you assume that the domain is all real numbers, you want to find the values that are not included in the domain. -2 and 1 are in the domain so you don't worry about them.

OpenStudy (anonymous):

Don't worry about them for the range right because i have -2,1,1,4

OpenStudy (anonymous):

im sorry -2,1,-1,4

hero (hero):

@kozy, hire a personal tutor for these because you need someone with a whiteboard who can explain it to you.

OpenStudy (anonymous):

okay thanks for your help

OpenStudy (ash2326):

Have you found the domain ?

OpenStudy (anonymous):

i think its -2 and 1 but idk if im right

OpenStudy (ash2326):

Write the denominator here?

OpenStudy (anonymous):

(x+1)(x-4)

OpenStudy (ash2326):

Equate it to 0, what would you get for x?

OpenStudy (anonymous):

you'll get -1 and 4

OpenStudy (ash2326):

These values of x, will make denominator 0 and f(x) will be undefined for these values. So domain is all real numbers except -1 and 4 \[R-\text{{-1, -4}}\]

OpenStudy (anonymous):

How about range that's the hard one for me

OpenStudy (ash2326):

Sorry, I made a mistake It'd be \[R-\text{{-1, 4}}\]

OpenStudy (ash2326):

Let's find the range :)

OpenStudy (ash2326):

\[F(x)=\frac{(x+2)(x-1)}{(x+1)(x-4)}\] Let f(x) be y \[y=\frac{(x+2)(x-1)}{(x+1)(x-4)}\] Can you find x in terms of y.??

OpenStudy (ash2326):

@kozy ???

hero (hero):

Proper Notation \[\left \{x \in R : x \ne-1, x \ne 4 \right \}\]

OpenStudy (anonymous):

I'm sorry what do you mean by finding x in terms of y?

OpenStudy (ash2326):

Suppose I have \[y=2x\] so x in terms of y is \[x=\frac{y}{2}\] You have to cross multiply and solve for x

OpenStudy (anonymous):

oh okay let me try

OpenStudy (ash2326):

take your time :)

OpenStudy (anonymous):

I'm sorry i'm stuck i don't know how to do that with the function we're using

OpenStudy (ash2326):

Oops I just realized that's it's difficult here, put x= infinity, what'd you get for y?

OpenStudy (anonymous):

infinity, 1 union infinity -4 ?

OpenStudy (ash2326):

I mean what you'd get for f(x), if you put x=\(\infty\) in f(x)

OpenStudy (anonymous):

i'm stuck :(

OpenStudy (ash2326):

ok, tell me what's \[\frac 1 \infty\]

OpenStudy (anonymous):

0

OpenStudy (ash2326):

good, divide numerator and denominator by x \[y=\frac{(x+2)(x-1)}{(x+1)(x-4)}\]

OpenStudy (ash2326):

then put x=infinity

OpenStudy (ash2326):

\[\large y=\frac{(1+\frac 2 x )(1-\frac 1x )}{(1+\frac 1 x)(1-\frac 4x)}\] now put x= \(\infty\)

OpenStudy (anonymous):

okay i got it! now which side is the range?

OpenStudy (ash2326):

what did you get for y?

hero (hero):

Hey @ash2326, help him with his previous problem.

OpenStudy (anonymous):

(1+2/infinity)(1-1/infinity) ? im a girl :)

hero (hero):

By the way, finding the range is much simpler than you think.

OpenStudy (ash2326):

\[\large y=\frac{(1+\frac 2 x )(1-\frac 1x )}{(1+\frac 1 x)(1-\frac 4x)}\] put x=\(\infty\) \[y=\frac{(1+0)(1-0)}{(1+0)(1-0)}=????\]

OpenStudy (anonymous):

oh right!! im sorry

OpenStudy (anonymous):

so y = 1?

OpenStudy (ash2326):

yes, this is the value of y when x is infinity. so range is all real no.s except 1

hero (hero):

Are you sure about that @ash2326?

OpenStudy (anonymous):

thank you

hero (hero):

I'm 100% sure the range is all real numbers

OpenStudy (anonymous):

are you sure @hero

hero (hero):

Yes. The reason why is because if you substitute f(x) with any number, you will be able to solve for x. If you are able to solve for x after substituting f(x) with a number, then that number is in the range.

hero (hero):

So if we take f(x) = 1 and we are able to solve for x, then 1 is in the range of the function. We need to find values that are out of the range. But we won't find any such number.